[figure: 9.7 cm]

TAILORING OF LOW-FIELD MAGNETORESISTANCE IN MANGANITE THIN FILMS

E.S. Vlakhov^{1,b}, K. Dörr^{a}, K.-H. Müller^{a}, A. Handstein^{a}, D. Eckert^{a}, K.A. Nenkov^{a,b}, T. Walter^{a}, R.A. Chakalov^{c}, R.I. Chakalova^{c}, A.Y. Spasov^{c}

1Institute of Solid State Physics, Bulgarian Academy of Sciences, Sofia, Bulgaria

aInstitut für Festkörper- und Werkstofforschung Dresden, Germany bInt. Laboratory for High Magnetic Fields and Low Temperatures, Wrocław, Poland

cInstitute of Electronics, Bulgarian Academy of Sciences, Sofia, Bulgaria

Magnetron sputtering and pulsed laser deposition techniques have been employed to deposit single crystal or polycrystalline manganite thin films on LaAlO_{3}(100), YSZ(100), MgO(100), Si(100)/YSZ, Si(111)/YSZ, Si(100)/MgO and Si(111)/MgO substrates. It was found that grain boundaries play an important role in low-field magnetoresistance (LFMR) of manganites. The influence of extrinsic properties of single-layer manganite thin films (La_{0.7}Ca_{0.3}MnO_{3} or La_{0.7}Sr_{0.3}MnO_{3}), modified by deposition conditions, lattice mismatch between the film and substrate/buffer layer, film thickness, etc. on LFMR sensitivity in wide temperature range up to Curie temperature (T_{C}) has been investigated. Well-pronounced LFMR hysteresis was obtained in magnetic fields applied in in-plane and out-of-plane direction. The high value of MR sensitivity d(MR)/dH in the vicinity of the MR peak (up to 225%/kOe at 5 K for samples on YSZ substrate) has to be noted. The obtained data have been analyzed using different theoretical models in order to tailor structure suitable for application.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : Dr. K. Dörr

Address for correspondence : IFW Dresden, Postfach 270016, 01171 Dresden, Germany

Email address : k.doerr@ifw-dresden.de